# Monomial web basis for the SL(N) skein algebra of the twice punctured sphere

### v1

@@ Acknowledgements
All vector spaces $V$ will be over $\mathbb{C}$. The notation $v\propto w$ means $v=aw$ for some scalar $a$. If $A$ is a, possibly non-commutative, algebra with $1$ over $\mathbb{C}$, denote by $\mathrm{sub}\left(a_{1},a_{2},\dots,a_{m}\right)$ the sub-algebra generated by elements $a_{1},a_{2},\dots,a_{m}\in A$, namely the linear span over $\mathbb{C}$ of all monomials of the form $a_{i_{1}}a_{i_{2}}\cdots a_{i_{p}}$ for $i_{j}\in\{1,2,\dots,m\}$, where this monomial is $1$ if $p=0$.
Let $\mathrm{Sym}_{m}$ denote the symmetric group on $m$ elements, namely $\mathrm{Sym}_{m}$ is the set of self bijections of $\{1,2,\dots,m\}$. We use the notation $\left(\begin{matrix}\sigma(1)&\sigma(2)&\ldots&\sigma(m)\end{matrix}\right)$ for permutations $\sigma\in\mathrm{Sym}_{m}$. For example, $\left(\begin{matrix}3&1&2&4\end{matrix}\right)\in\mathrm{Sym}_{4}$ denotes the permutation where $1\mapsto 3$, $2\mapsto 1$, $3\mapsto 2$, $4\mapsto 4$. The order of multiplication in $\mathrm{Sym}_{m}$ coincides with the usual convention for multiplying functions, so if $\sigma_{1},\sigma_{2}\in\mathrm{Sym}_{m}$, then the composition $\sigma_{2}\circ\sigma_{1}\in\mathrm{Sym}_{m}$ is defined by $(\sigma_{2}\circ\sigma_{1})(i)=\sigma_{2}(\sigma_{1}(i))$ for $i=1,2,\dots,m$. The length $\ell(\sigma)$ of a permutation is the minimum number of factors appearing in a decomposition of $\sigma$ as the product of adjacent transpositions.
Curves in two- and three-dimensional spaces are always oriented. When permutations $\sigma\in\mathrm{Sym}_{m}$ are represented by overlapping upward oriented curves in a box, they are read from bottom to top. For example, the picture on the left hand side of Figure 1 represents the permutation $\left(\begin{matrix}3&1&2&4\end{matrix}\right)\in\mathrm{Sym}_{4}$.
Figure 1. The permutation $\left(\begin{matrix}3&1&2&4\end{matrix}\right)$ and its corresponding positive permutation braid.
@@
We use the notation $\widehat{\beta}\in\mathrm{Sym}_{m}$ to denote the permutation projected to by the braid $\beta$. The multi-curve $\widehat{\beta}$ has no monogons, but, before any homotopy, could have bigons. Here, by monogon and bigon, we mean as in Figure 2, where on the left hand side the contractible enclosed region is a monogon, and where on the right hand side the contractible enclosed region is a bigon. Note that it is possible for this enclosed region to intersect other parts of the multi-curve. As an example of this subtlety about homotopy, the braid $\beta\in\mathrm{Brd}_{2}$ representing a full twist satisfies that $\widehat{\beta}=\left(\begin{matrix}2&1\end{matrix}\right)^{2}$ has a bigon, even though as a permutation it is trivial after homotopy. The permutation picture $\widehat{\beta}$ can always be homotoped into minimal position, that is, so that it has no bigons. We call $\beta$ a permutation braid if $\beta$ can be isotoped such that $\widehat{\beta}$ has no bigons. Every permutation $\sigma$ has $2^{\ell(\sigma)}$ lifts, denoted by $\widetilde{\sigma}$, which are just the permutation braids projecting to $\sigma$. For example, $\widetilde{\sigma}_{+}$ and $\widetilde{\sigma}_{-}$ are two such lifts. Note that the permutation braids lifting the same permutation are just the mutants of $\widetilde{\sigma}_{+}$.
@@ 1.2. Skeins
The skein algebra $\mathcal{S}_{n}^{q}$ is denoted in the notation of [46] by $\mathcal{S}_{n}^{b}(M;u)$ for $M=\mathfrak{A}\times\mathfrak{I}$ and $u=(-q)^{n(n-1)/2}$, and was first defined, for general $n$, in [66]. The unbased skein algebra $\left(\mathcal{S}_{n}^{q}\right)^{\circ}$ was first defined, for general $n$, in [46] and is denoted by $\mathcal{S}_{n}(M,\varnothing)$ in the notation of [46].
@@
For all $q^{1/n}\in\mathbb{C}^{\ast}$, a collection $\{W_{i}\}_{i\in I}$ of webs forms a linear basis for $\mathcal{S}_{n}^{q}$ if and only if the collection of underlying unbased webs $\{W_{i}^{\circ}\}_{i\in I}$ forms a linear basis for $\left(\mathcal{S}_{n}^{q}\right)^{\circ}$. ∎
@@ 3.1. Setup
For a permutation $\sigma\in\mathrm{Sym}_{m}$, the permutation multi-curve associated to $\sigma$, which by abuse of notation we also denote by $\sigma$, is the oriented multi-curve in $\mathfrak{A}$ obtained from the permutation picture for $\sigma$ in a box in $\mathfrak{A}$ by connecting the boundary strands with parallel curves going counterclockwise around the annulus. See for example the leftmost picture in Figure 7 which is the permutation multi-curve $\sigma=\left(\begin{matrix}3&1&2&4\end{matrix}\right)$. Permutation multi-curves are considered up to homotopy in the annulus. Note that the composition $\sigma_{2}\circ\sigma_{1}$ of permutations makes sense as a permutation multi-curve.
@@
Similarly, for a braid $\beta\in\mathrm{Brd}_{m}$ we can consider the braid link associated to $\beta$, which by abuse of notation we also denote by $\beta$, as the link in $\mathfrak{A}\times\mathfrak{I}$ obtained by closing up the boxed braid counterclockwise around the thickened annulus. See for example the middle and rightmost pictures in Figure 7. Note that the rightmost braid is the positive braid $\widetilde{\sigma}_{+}$ for $\sigma=\left(\begin{matrix}2&3&4&1\end{matrix}\right)$. Braid links are considered up to isotopy in the thickened annulus. Note that the composition $\beta_{2}\circ\beta_{1}$ of braids makes sense as a braid link. We will also think of a braid link $\beta$ as an element of $\mathcal{S}_{n}^{q}$. Note, in particular, that the composition $\beta_{2}\circ\beta_{1}$ is different from the product $\beta_{2}\beta_{1}$ in the skein algebra $\mathcal{S}_{n}^{q}$. The reverse braid link $\overline{\beta}$ is obtained by reversing the orientation of the braid link $\beta$. We refer to a connected braid link $\beta$ as a braid knot. A permutation link is a braid link associated to a permutation braid $\beta=\widetilde{\sigma}$ for some $\sigma\in\mathrm{Sym}_{m}$. A permutation knot is a connected permutation link.
@@ 3.3. Countable generating set
For $m\in\mathbb{Z}$, define the $m$-th power knot $\gamma_{m}$ in $\mathfrak{A}\times\mathfrak{I}$ as follows. For $m>0$, we put $\gamma_{m}$ to be the permutation knot $\gamma_{m}=\widetilde{\left(\begin{matrix}2&3&\ldots&m&1\end{matrix}\right)}_{+}$. For $m<0$, $\gamma_{m}$ is the reverse permutation knot $\overline{\gamma_{|m|}}$. And $\gamma_{0}=\varnothing$. For example, the rightmost picture in Figure 7 is $\gamma_{4}$. We will also think of $\gamma_{m}$ as an element of $\mathcal{S}_{n}^{q}$. Note that in the fundamental group of the annulus, the projection $\widehat{\gamma_{m}}$ represents the $m$-th power of the counterclockwise generator. Then, we have:
@@
  1. (1)
    It suffices to show that $\left[W\right]\in\mathcal{G}$ for any web $W$. Indeed, every element of $\mathcal{S}_{n}^{q}$ can be written as a linear combination of classes of webs.
  2. (2)
    It suffices to show that $\left[L\right]\in\mathcal{G}$ for any link $L$. Indeed, by iteratively applying the vertex removing relation, the class $\left[W\right]$ of a web can be written as a linear combination of classes of links.
  3. (3)
@@ 3.4. Main result of the section for the bad set: statement
($\ast$) $B_{i}-q^{n(n-1)}\left(\prod_{k=i}^{n-1}q^{-k}[n-k]\right)P_{i}(q^{1/n})\gamma_{i}\in\mathrm{sub}\left(\gamma_{1},\gamma_{2},\dots,\gamma_{i-1}\right)$
and
@@ 3.5. Main result of the section for the bad set: induction step
Let $a=q^{n(n-1)}$, $b=-q^{(1-n)/n}$, and $a_{k}=q^{-k}[n-k]$. Then, for all $q^{1/n}\in\mathbb{C}^{\ast}$, the relation appearing in Figure 10 holds. The same relation holds with all of the orientations reversed.
@@ 3.8. Digression: countable generating set for general surfaces
In this sub-section, we digress to prove a generalization of Proposition 3.2, which will not be used in the rest of the paper. This is a very slight generalization of a well-known result for the HOMFLYPT skein algebra [68] of a surface. For completeness, we give a proof in the arXiv version of this article.
@@
For an oriented surface $\mathfrak{S}$, we have $\mathcal{S}_{n}^{q}=\mathrm{sub}\left(\mathcal{G}_{\alpha};\alpha\in\Pi\right)$.
@@ 4.1. Stated webs
In order to represent a stated web, rather, its isotopy class, by a diagram, we need to indicate in the diagram the height data associated to points $W^{\circ}\cap(\mathfrak{K}\times\mathfrak{I})$ of the web in each boundary wall. This can be achieved by drawing a white arrow on the boundary component $\mathfrak{K}$, together with a starting location, such as in Figure 14, where the heights of the boundary points increase as one moves in the direction of the arrow. As an example, in relation (B) of Figure 14, the point with state $j$ is higher than the point with state $i$, where in the figure we have assumed the starting location for the arrow on the boundary component $\mathfrak{K}$ is to the right. Note that such diagrams always exist, by horizontally sliding via isotopy the boundary points of $W^{\circ}$ along a boundary wall. Note also that such an arrow is only necessary when $W^{\circ}\cap(\mathfrak{K}\times\mathfrak{I})$ consists of at least two points. Soon, we will work exclusively with punctured surfaces (Section 4.3), where the starting location is determined by the arrow and is therefore superfluous.
@@ 4.3. Punctured surfaces and the biangle quantum trace map
The stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\mathfrak{D}_{1})$ of the monoangle is isomorphic as an algebra to the complex numbers $\mathbb{C}$. The stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\mathfrak{D}_{2})$ of the biangle is a Hopf algebra isomorphic to the quantized coordinate ring $\mathcal{O}_{q}(\mathrm{SL}(n))$. The stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\mathfrak{D}_{3})$ of the triangle is isomorphic to the braided tensor product of $\mathcal{O}_{q}(\mathrm{SL}(n))$ with itself.
@@
For $k\in\{0,1,\dots,n\}$ and $i_{j}\in\{1,2,\dots,n\}$ for $j=1,2,\dots,k$ and $l_{j^{\prime}}\in\{1,2,\dots,n\}$ for $j^{\prime}=1,2,\dots,n-k$, let $W^{\circ}_{\mathrm{source}}(k,(i_{j})_{j},(l_{j^{\prime}})_{j^{\prime}})$ and $W^{\circ}_{\mathrm{sink}}(k,(i_{j})_{j},(l_{j^{\prime}})_{j^{\prime}})$ be the source and sink stated webs in the biangle $\mathfrak{D}_{2}$ appearing in Figure 16. The co-unit $\epsilon$ has the same values on these two stated webs, which we provide in the figure. Here, if $\{i_{1},i_{2},\dots,i_{k},\overline{l}_{n-k},\dots,\overline{l}_{2},\overline{l}_{1}\}$ has less than $n$ elements then the co-unit is zero, else $\left(\begin{matrix}i_{1}&i_{2}&\ldots&i_{k}&\overline{l}_{n-k}&\ldots&\overline{l}_{2}&\overline{l}_{1}\end{matrix}\right)\in\mathrm{Sym}_{n}$ is the permutation determined by the boundary states in this order.
@@
which is equal to the right hand side by the general formula $\sum_{\sigma\in\mathrm{Sym}_{k}}(-q)^{2\ell(\sigma)}=q^{k(k-1)/2}[k]!$; see, for example, [66, Equation (21)]. ∎

### v2

@@ Acknowledgements
Vector spaces $V$ will be over $\mathbb{C}$. The notation $v\propto w$ means $v=aw$ for some scalar $a$. If $A$ is a, possibly non-commutative, algebra with $1$ over $\mathbb{C}$, denote by $\mathrm{sub}\left(a_{1},a_{2},\dots,a_{m}\right)$ the sub-algebra generated by elements $a_{1},a_{2},\dots,a_{m}\in A$, namely the linear span over $\mathbb{C}$ of all monomials of the form $a_{i_{1}}a_{i_{2}}\cdots a_{i_{p}}$ for $i_{j}\in\{1,2,\dots,m\}$, where this monomial is $1$ if $p=0$.
Let $\mathrm{Sym}_{m}$ denote the symmetric group on $m$ elements, so $\mathrm{Sym}_{m}$ is the set of self bijections of $\{1,2,\dots,m\}$. We denote permutations $\sigma\in\mathrm{Sym}_{m}$ by $\left(\begin{smallmatrix}\sigma(1)&\sigma(2)&\ldots&\sigma(m)\end{smallmatrix}\right)$. For example, $\left(\begin{smallmatrix}3&1&2&4\end{smallmatrix}\right)\in\mathrm{Sym}_{4}$ denotes the permutation where $1\mapsto 3$, $2\mapsto 1$, $3\mapsto 2$, $4\mapsto 4$. The order of multiplication in $\mathrm{Sym}_{m}$ coincides with the usual convention for multiplying functions, so if $\sigma_{1},\sigma_{2}\in\mathrm{Sym}_{m}$, then the composition $\sigma_{2}\circ\sigma_{1}\in\mathrm{Sym}_{m}$ is defined by $(\sigma_{2}\circ\sigma_{1})(i)=\sigma_{2}(\sigma_{1}(i))$ for $i=1,2,\dots,m$. The length $\ell(\sigma)$ of a permutation is the minimum number of factors appearing in a decomposition of $\sigma$ as the product of adjacent transpositions.
Curves in two- and three-dimensional spaces are always oriented. When permutations $\sigma\in\mathrm{Sym}_{m}$ are represented by overlapping upward oriented curves in a box, they are read from bottom to top. For example, the picture on the left hand side of Figure 1 represents the permutation $\left(\begin{smallmatrix}3&1&2&4\end{smallmatrix}\right)\in\mathrm{Sym}_{4}$.
Figure 1. The permutation $\left(\begin{smallmatrix}3&1&2&4\end{smallmatrix}\right)$ and its corresponding positive permutation braid.
@@
We use the notation $\widehat{\beta}\in\mathrm{Sym}_{m}$ to denote the permutation projected to by the braid $\beta$. The multi-curve $\widehat{\beta}$ has no monogons, but, before any homotopy, could have bigons. Here, by monogon and bigon, we mean as in Figure 2, where on the left hand side the contractible enclosed region is a monogon, and where on the right hand side the contractible enclosed region is a bigon. Note that it is possible for this enclosed region to intersect other parts of the multi-curve. As an example of this subtlety about homotopy, the braid $\beta\in\mathrm{Brd}_{2}$ representing a full twist satisfies that $\widehat{\beta}=\left(\begin{smallmatrix}2&1\end{smallmatrix}\right)^{2}$ has a bigon, even though as a permutation it is trivial after homotopy. The permutation picture $\widehat{\beta}$ can always be homotoped into minimal position, that is, so that it has no bigons. We call $\beta$ a permutation braid if $\beta$ can be isotoped such that $\widehat{\beta}$ has no bigons. Every permutation $\sigma$ has $2^{\ell(\sigma)}$ lifts, denoted by $\widetilde{\sigma}$, which are just the permutation braids projecting to $\sigma$. For example, $\widetilde{\sigma}_{+}$ and $\widetilde{\sigma}_{-}$ are two such lifts. Note that the permutation braids lifting the same permutation are just the mutants of $\widetilde{\sigma}_{+}$.
@@ 1.2. Skeins
The skein algebra $\mathcal{S}_{n}^{q}$ is denoted in the notation of [49] by $\mathcal{S}_{n}^{b}(M;u)$ for $M=\mathfrak{A}\times\mathfrak{I}$ and $u=(-q)^{n(n-1)/2}$, and was first defined, for general $n$, in [69]. The unbased skein algebra $\left(\mathcal{S}_{n}^{q}\right)^{\circ}$ was first defined, for general $n$, in [49] and is denoted by $\mathcal{S}_{n}(M,\varnothing)$ in [49].
@@
For all $q^{1/n}\in\mathbb{C}^{\ast}$, a collection $\{W_{i}\}_{i\in I}$ of webs forms a linear basis for $\mathcal{S}_{n}^{q}$ if and only if the collection of underlying unbased webs $\{W_{i}^{\circ}\}_{i\in I}$ forms a basis for $\left(\mathcal{S}_{n}^{q}\right)^{\circ}$. ∎
@@ 3.1. Setup
For a permutation $\sigma\in\mathrm{Sym}_{m}$, the permutation multi-curve associated to $\sigma$, which by abuse of notation we also denote by $\sigma$, is the oriented multi-curve in $\mathfrak{A}$ obtained from the permutation picture for $\sigma$ in a box in $\mathfrak{A}$ by connecting the boundary strands with parallel curves going counterclockwise around the annulus. See for example the leftmost picture in Figure 7 which is the permutation multi-curve $\sigma=\left(\begin{smallmatrix}3&1&2&4\end{smallmatrix}\right)$. Permutation multi-curves are considered up to homotopy in the annulus. Note that the composition $\sigma_{2}\circ\sigma_{1}$ of permutations makes sense as a permutation multi-curve.
@@
Similarly, for a braid $\beta\in\mathrm{Brd}_{m}$ we can consider the braid link associated to $\beta$, which by abuse of notation we also denote by $\beta$, as the link in $\mathfrak{A}\times\mathfrak{I}$ obtained by closing up the boxed braid counterclockwise around the thickened annulus. See for example the middle and rightmost pictures in Figure 7. Note that the rightmost braid is the positive braid $\widetilde{\sigma}_{+}$ for $\sigma=\left(\begin{smallmatrix}2&3&4&1\end{smallmatrix}\right)$. Braid links are considered up to isotopy in the thickened annulus. Note that the composition $\beta_{2}\circ\beta_{1}$ of braids makes sense as a braid link. We will also think of a braid link $\beta$ as an element of $\mathcal{S}_{n}^{q}$. Note, in particular, that the composition $\beta_{2}\circ\beta_{1}$ is different from the product $\beta_{2}\beta_{1}$ in the skein algebra $\mathcal{S}_{n}^{q}$. The reverse braid link $\overline{\beta}$ is obtained by reversing the orientation of the braid link $\beta$. We refer to a connected braid link $\beta$ as a braid knot. A permutation link is a braid link associated to a permutation braid $\beta=\widetilde{\sigma}$ for some $\sigma\in\mathrm{Sym}_{m}$. A permutation knot is a connected permutation link.
@@ 3.3. Countable generating set
For $m\in\mathbb{Z}$, define the $m$-th power knot $\gamma_{m}$ in $\mathfrak{A}\times\mathfrak{I}$ as follows. For $m>0$, we put $\gamma_{m}$ to be the permutation knot $\gamma_{m}=\widetilde{\left(\begin{smallmatrix}2&3&\ldots&m&1\end{smallmatrix}\right)}_{+}$. For $m<0$, $\gamma_{m}$ is the reverse permutation knot $\overline{\gamma_{|m|}}$. And $\gamma_{0}=\varnothing$. For example, the rightmost picture in Figure 7 is $\gamma_{4}$. We will also think of $\gamma_{m}$ as an element of $\mathcal{S}_{n}^{q}$. Note that in the fundamental group of the annulus, the projection $\widehat{\gamma_{m}}$ represents the $m$-th power of the counterclockwise generator. Then, we have:
@@
(1) It suffices to show that $\left[W\right]\in\mathcal{G}$ for any web $W$. Indeed, every element of $\mathcal{S}_{n}^{q}$ can be written as a linear combination of classes of webs.
(2) It suffices to show that $\left[L\right]\in\mathcal{G}$ for any link $L$. Indeed, by iteratively applying the vertex removing relation, the class $\left[W\right]$ of a web can be written as a linear combination of classes of links.
@@ 3.4. Main result of the section for the bad set: statement
($\ast$) $B_{i}-q^{n(n-1)}\left(\prod_{k=i}^{n-1}q^{-k}[n-k]\right)P_{i}(q^{1/n})\gamma_{i}\in\mathrm{sub}\left(\gamma_{1},\gamma_{2},\dots,\gamma_{i-1}\right),$
@@ 3.5. Main result of the section for the bad set: induction step
Let $a=q^{n(n-1)}$, $b=-q^{(1-n)/n}$, and $a_{k}=q^{-k}[n-k]$. For all $q^{1/n}\in\mathbb{C}^{\ast}$, the relation appearing in Figure 10 holds. The same relation holds with all orientations reversed.
@@ 3.8. Digression: countable generating set for general surfaces
In this sub-section, we digress to prove a generalization of Proposition 3.2, which will not be used in the rest of the paper. This is a very slight generalization of a well-known result for the HOMFLYPT skein algebra [73] of a surface.
@@
For an oriented surface $\mathfrak{S}$, we have $\mathcal{S}_{n}^{q}=\mathrm{sub}\left(\mathcal{G}_{\alpha};\alpha\in\Pi\right)$. ∎
@@ 4.1. Stated webs
In order to represent a stated web, rather, its isotopy class, by a diagram, we need to indicate in the diagram the height data associated to points $W^{\circ}\cap(\mathfrak{K}\times\mathfrak{I})$ of the web in each boundary wall. This can be achieved by drawing a white arrow on the boundary component $\mathfrak{K}$, together with a starting location, such as in Figure 14, where the heights of the boundary points increase as one moves in the direction of the arrow. As an example, in relation (B) of Figure 14, the point with state $j$ is higher than the point with state $i$, where in the figure we have assumed the starting location for the arrow on the boundary component $\mathfrak{K}$ is to the right. Note that such diagrams always exist, by horizontally sliding via isotopy the boundary points of $W^{\circ}$ along a boundary wall. Note also that such an arrow is only necessary when $W^{\circ}\cap(\mathfrak{K}\times\mathfrak{I})$ consists of at least two points. Soon, we will work only with punctured surfaces (Section 4.3), where the starting location is determined by the arrow so is superfluous.
@@ 4.3. Punctured surfaces and the biangle quantum trace map
The stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\mathfrak{D}_{1})$ of the monoangle is isomorphic as an algebra to $\mathbb{C}$. The stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\mathfrak{D}_{2})$ of the biangle is a Hopf algebra isomorphic to the quantized coordinate ring $\mathcal{O}_{q}(\mathrm{SL}(n))$. The stated skein algebra $\widetilde{\mathcal{S}}_{n}^{q}(\mathfrak{D}_{3})$ of the triangle is isomorphic to the braided tensor product of $\mathcal{O}_{q}(\mathrm{SL}(n))$ with itself.
@@
For $k\in\{0,1,\dots,n\}$ and $i_{j}\in\{1,2,\dots,n\}$ for $j=1,2,\dots,k$ and $l_{j^{\prime}}\in\{1,2,\dots,n\}$ for $j^{\prime}=1,2,\dots,n-k$, let $W^{\circ}_{\mathrm{source}}(k,(i_{j})_{j},(l_{j^{\prime}})_{j^{\prime}})$ and $W^{\circ}_{\mathrm{sink}}(k,(i_{j})_{j},(l_{j^{\prime}})_{j^{\prime}})$ be the source and sink stated webs in the biangle $\mathfrak{D}_{2}$ appearing in Figure 16. The co-unit $\epsilon$ has the same values on these two stated webs, which we provide in the figure. Here, if $\{i_{1},i_{2},\dots,i_{k},\overline{l}_{n-k},\dots,\overline{l}_{2},\overline{l}_{1}\}$ has less than $n$ elements then the co-unit is zero, else $\left(\begin{smallmatrix}i_{1}&i_{2}&\ldots&i_{k}&\overline{l}_{n-k}&\ldots&\overline{l}_{2}&\overline{l}_{1}\end{smallmatrix}\right)\in\mathrm{Sym}_{n}$ is the permutation determined by the boundary states in this order.
@@
which is equal to the right hand side by the general formula $\sum_{\sigma\in\mathrm{Sym}_{k}}(-q)^{2\ell(\sigma)}=q^{k(k-1)/2}[k]!$; see, for example, [69, Equation (21)]. ∎